It is 2800 miles from new York to London.it takes 7 hours to get to London with the wind and 9 hours on a return trip against the wind. What is the rate of the wind
Answer by TimothyLamb(4379)

s = d / t
t = d / s
---
2800/(a + w) = 7
2800 = 7(a + w)
2800 = 7a + 7w
7a = 2800 - 7w
a = 400 - w
---
2800/(a - w) = 9
2800 = 9(a - w)
2800 = 9a - 9w
---
2800 = 9a - 9w
2800 = 9(400 - w) - 9w
2800 = 3600 - 9w - 9w
18w = 800
---
wind speed = 44.444444 mph
airplane speed = 355.555555 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php